# Human Creativity in the Age of LLMs

### v1

@@ Abstract.
conference: ; September 2024; Toronto, ON, Canadaccs: Human-centered computing Empirical studies in HCIccs: Applied computing Arts and humanitiesccs: Human-centered computing Laboratory experiments
@@ 2.2. LLMs as Tools for Creativity
Large Language Models (LLMs) have demonstrated notable creative performance, often performing as well as, or even better than, average humans in various creative tasks. However, they still fall short when compared to the best human performers, particularly experts (Koivisto and Grassini, 2023; Hubert et al., 2024). For instance, Chakrabarty et al. (Chakrabarty et al., 2024) found that while LLMs performed competently in creative writing tasks, professionals consistently outperformed them. Anderson et al. (Anderson et al., 2024) observed that ChatGPT users generated more ideas than those in a control condition. However, these ideas tended to be homogenized across different users, suggesting a limitation in the diversity of creative output when LLMs are used. These findings are particularly relevant when LLMs are employed to perform entire tasks on behalf of users.
@@ 2.3. Generative AI and Human Cognition
Although the long-term effects of Generative AI use has been studied empirically in domains such as education (for e.g., (Kumar et al., 2023; Bastani et al., 2024)), web search (for e.g., (Spatharioti et al., 2023)), etc. There is a lack of empirical evidence on the effect of Generative AI tools on our convergent and divergent thinking abilities.
@@ 3. Experiment 1: Divergent Thinking
A critical aspect of creativity is the ability to generate a wide range of high-quality ideas, often called divergent thinking (Runco, 1991; Guilford, 1967). LLMs have the capacity to generate a significant quantity of ideas, often exceeding what an individual can produce, without the constraints of time or context. However, the impact of using LLM-generated ideas on an individual’s ability to think divergently and come up with ideas independently remains underexplored. In addition to producing ideas, LLMs can offer structured frameworks that guide users in their creative processes, much like a coach (Baer, 2014). To understand these dynamics, our first pre-registered [^1] https://aspredicted.org/4X6_6X6 experiment investigates the effects of LLMs that either directly provide ideas or guide participants using a framework, helping them to develop their own ideas.
Figure 2. Schematic of design for Experiment 1 on divergent thinking.
@@ 3.1.1. Experimental conditions and treatment
Our experimental design involved two main phases: an exposure (paired) phase during which participants attempted the task with an LLM partner, and a test (solo) phase during which participants attempted the task on their own. Participants were given two-minute time frames per object to submit their ideas one by one, and could freely edit or delete any previously submitted ideas. During the exposure phase, participants are introduced to three objects, one after the other. Every participant was randomly assigned to receive one of three types of LLM responses:
@@
The assigned type of LLM response appeared 5 seconds after the item was shown to the participant, and appeared character by character, similar to other chat LLM interfaces. Figure 2 shows the schema of the experiment design. Following the exposure phase, participants engaged in a brief distractor task to simulate forgetting, playing a game of Snake for one minute. In this subsequent phase, participants were assigned the Alternate Uses Task for a new object selected at random, this time without LLM support.
@@ 3.2.3. Creative Flexibility
Figure 5. Plots of the average individual- and group-level median diversity, segmented by experiment conditions and phases. Higher values denote more difference between ideas. Error bars represent $\pm$ one standard error of mean.
@@ 4. Experiment 2: Convergent Thinking
The results from our first experiment highlight how LLMs shape our ability to generate ideas independently. However, creativity involves more than idea generation; it requires the capacity to identify and refine the most appropriate solutions within a given conceptual space. Convergent thinking is crucial for advancing from a broad set of ideas to selecting the most viable outcome. Motivated by this, our second pre-registered [^2] https://aspredicted.org/V8K_ZTV experiment focuses on convergent thinking, testing not only LLM-generated answers but also the potential of LLMs to guide users toward solutions, simulating a coaching approach.
@@ 4.1.1. Conditions and treatment
After the exposure phase, participants played a game of Snake for one minute which served as a distractor task. In the test phase, all participants completed two additional RAT rounds (randomly selected) to attempt without any LLM assistance.
@@
During the exposure rounds, participants were made to wait the full 1 minute before advancing to the next task, while in the test rounds, participants could advance once they submitted an answer. This was to discourage participants from speeding through the experiment without giving thought to their answers. Participants were never shown the correct answer to the RATs they completed (except as part of LLM responses, if applicable), nor were they informed if their answers were correct.
@@ 4.1.2. Stimulus
For the Remote Associates Test (RAT), we employed a widely-used dataset from Bowden and Jung-Beeman (Bowden and Jung-Beeman, 2003). The subset we selected contained 45 questions equally distributed across three difficulty levels: easy, medium, and hard. For our study, we randomly selected questions from the easy and medium difficulty levels for each round, as pilot testing indicated that these were appropriate for our participant pool (crowdworkers). Figure 14 provides examples of the LLM responses presented to participants in the different experimental conditions.
@@ 5. Discussion
In our discussion, we first summarize the key findings of our work and emphasize our contribution as it relates to creativity and the evaluation of AI-based support tools. We then highlight our works’ contributions to the broader implications on fields involving co-creativity with AI. Finally, we discuss the limitations of our study and avenues for future research.
@@ 5.1.2. Differential Impact of LLMs on Divergent and Convergent Thinking
Our experiments reveal that the effects of LLMs vary significantly depending on the aspect of creativity being measured. In divergent thinking tasks, where participants were asked to generate a wide range of ideas, we observed more skepticism toward LLM assistance. Participants seemed less inclined to adopt AI-generated suggestions, which may be due to the nature of divergent thinking itself—encouraging exploration and unconventional approaches (Guilford, 1967; Runco, 1991). In contrast, for the convergent thinking task, where participants were tasked with narrowing down the ideas to a single solution, LLM assistance during the exposure phase appeared more beneficial. This is consistent with the idea that convergent thinking is more structured and goal-oriented, making it easier for participants to recognize when LLMs are effectively guiding them toward the correct solution (Mednick and Halpern, 1968).
LLMs, however, may also hinder the creative process. In divergent thinking, the introduction of AI-generated ideas may distract participants, consume valuable cognitive resources, or prevent full engagement with the task. This aligns with theories of creativity such as Boden’s concept of conceptual spaces, which emphasize the importance of users exploring and navigating creative possibilities independently (Boden, 2010). Over-reliance on LLMs during divergent thinking may disrupt this exploration, leading to less original ideas, as seen in the originality results. Similarly, in convergent thinking tasks, LLMs might steer participants toward specific solutions too quickly, reducing the need for deep engagement with the problem space (Guilford, 1967).
@@ 5.1.3. Persistent Homogenization of Ideas and the Challenges of Designing LLM Coaches
Existing research has shown that LLM usage can lead to the homogenization of ideas within groups, where participants tend to converge on similar outcomes when using AI-generated suggestions (Anderson et al., 2024). Our findings extend this concern, showing that even when people stop using LLMs that provide strategic frameworks for thinking, the homogenization effect can persist. In our divergent thinking experiment, participants who received LLM-generated strategies exhibited reduced diversity in their idea sets both during and after LLM use, suggesting that such frameworks may have a lasting impact on creative processes, potentially stifling the generation of more varied or unconventional ideas.
Interestingly, we did not find the same effect for participants who received direct ideas from a standard LLM. The List of Ideas condition did not lead to the same lasting homogenization once the LLM was no longer present. This indicates that while LLM-generated strategies can have long-term effects on creative diversity, providing ideas without a guiding framework may allow for more cognitive flexibility once participants stop using AI. In our convergent thinking experiment, participants who received LLM guidance during exposure performed worse in the test round compared to those who received direct answers. This highlights the complexity of designing LLMs as coaches or guides. While direct answers may not always seem ideal, in some cases, they can be more effective than offering frameworks or strategies. Designers of Human-AI systems must take these findings into account, ensuring that LLM interactions are structured to avoid unintentional long-term effects on cognitive diversity. Controlled experiments, such as those conducted here, can be useful methods in refining and optimizing the design of coach-like LLM systems.
@@ 5.2. Broader Implications for Fields Involving Human-AI Co-Creativity
These insights have broad implications for various fields, from scientific discovery to the arts and humanities. In the context of AI for science, for instance, there is ongoing excitement about building increasingly sophisticated models to accelerate the process of hypothesis generation and experimental execution (Boiko et al., 2023; Si et al., 2024; Lu et al., 2024). However, the design of these models often overlooks their potential impact on scientists’ creative abilities. Although some initial explorations have been conducted, the field lacks rigorous empirical evaluation of how AI systems affect human creativity in scientific discovery. The arts and humanities may face similar challenges (Rane and Choudhary, 2024). For example, writers using the same LLM could produce homogenized content, even when AI is no longer part of their workflow. This underscores the need for AI systems that not only assist in the creative process but also promote long-term cognitive diversity, ensuring that human creativity thrives in collaboration with AI rather than becoming constrained by it.
Our work in a controlled, simple setting suggests that the performance of AI models alone is not enough to determine their value. The design of human-AI interactions—whether AI provides direct answers or encourages users to think critically—and the timing of performance evaluation (during AI use versus after) both significantly alter the narrative. These considerations are critical in ensuring that AI models enhance, rather than diminish, human creative potential. Beyond simply building superhuman AI, we must focus on how these tools influence human creativity, culture, and cognitive growth, aiming for AI systems that enrich and elevate human thought.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714198isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in HCIccs: Applied computing Arts and humanitiesccs: Human-centered computing Laboratory experiments
@@
Figure 1 depicts a pictorial flow chart illustrating the high level structural overview of the experiments performed. Participants are randomized into one of three conditions for three exposure rounds: A (human-only), B (AI provides full solutions), and C (AI guides but doesn’t provide full solutions). In Condition A, participants complete 3 tasks independently. In Condition B, participants complete 3 tasks with AI giving the solution, while in Condition C, participants complete 3 tasks with AI guidance without providing the final answer. In the final test round (D), all participants work independently, regardless of their previous condition.
@@ 2.2. LLMs as Tools for Creativity
LLMs have demonstrated notable creative performance, often performing as well as, or even better than, average humans in various creative tasks. However, they still fall short when compared to the best human performers, particularly experts (Koivisto and Grassini, 2023; Hubert et al., 2024). For instance, Chakrabarty et al. (Chakrabarty et al., 2024) found that while LLMs performed competently in creative writing tasks, professionals consistently outperformed them. Anderson et al. (Anderson et al., 2024) observed that ChatGPT users generated more ideas than those in a control condition. However, these ideas tended to be homogenized across different users, suggesting a limitation in the diversity of creative output when LLMs are used. Koivisto et al. (Koivisto and Grassini, 2023) compared LLM vs. human performance in a divergent thinking task, but without examining how LLM use affects human creativity. Ashkinaze et al. (Ashkinaze et al., 2024) conducted a single round idea generation with LLM support and asked participants to provide a single original idea. These studies are particularly relevant when LLMs are employed to perform entire tasks on behalf of users.
@@
Figure 2. Schematic of design for Experiment 1 on divergent thinking.
[…]
Figure 2 illustrates the high level structural overview of the divergent thinking experiment. 460 participants start with a consent form and pre-experiment survey before being randomized into one of three conditions for three exposure rounds. These conditions are: A (human-only), B (AI provides a list of ideas), and C (AI provides a list of brainstorming strategies). The items provided in these exposure rounds are randomly sampled from an item bank. After the exposure rounds, participants engage in a 1 minute snake game as a brief interlude before moving onto a test round that all participants carry out without any AI support. Finally, they end with a post-survey.
@@ 2.3. Generative AI and Human Cognition
Situating work in broader HCI literature.
Although the long-term effects of Generative AI use has been studied empirically in domains such as education (Kumar et al., 2023; Bastani et al., 2024), web search (Spatharioti et al., 2023), etc., there is a lack of empirical evidence on the effect of Generative AI tools on our convergent and divergent thinking abilities. Our approach differs significantly from previous work that focuses solely on measuring and improving the creative output of LLMs themselves (Lu et al., 2024a), without considering how their use affects human (user) creativity. Unlike previous work, we examine and discuss both aspects of creativity (divergent and convergent thinking) together for a comprehensive understanding of how LLMs shape human creative processes. Prior studies have used AUT and RAT to evaluate the creative abilities of LLMs (Lu et al., 2024a; Koivisto and Grassini, 2023) or to assess Human+AI creativity under a single exposure (Ashkinaze et al., 2024). In contrast, our work approximates real-world interactions by incorporating multiple sequential exposures to LLMs. This allowed us to capture how users adapt their creative processes over time, providing a more realistic assessment of the impacts of LLM-assisted creativity. In addition, we experimentally measure the impact of these repeated exposures on human creativity (residual effects of using LLMs), an aspect missing from existing research on LLM and creativity.
@@ 3. Experiment 1: Divergent Thinking
A critical aspect of creativity is the ability to generate a wide range of high-quality ideas, often called divergent thinking (Runco, 1991; Guilford, 1967). LLMs have the capacity to generate a significant quantity of ideas, often exceeding what an individual can produce, without the constraints of time or context. However, the impact of using LLM-generated ideas on an individual’s ability to think divergently and come up with ideas independently remains underexplored. In addition to producing ideas, LLMs can offer structured frameworks that guide users in their creative processes, much like a coach (Baer, 2014). To understand these dynamics, our first pre-registered [^1] https://aspredicted.org/dkvp-n785.pdf experiment investigates the effects of LLMs that either directly provide ideas or guide participants using a framework, helping them to develop their own ideas.
@@ 3.1. Experimental Design
[…]
Figure 3 shows the 4 user interfaces used in the divergent thinking experiment, (No LLM Response, LLM Answer, and LLM Guidance exposure conditions, as well as for the test round). All screens are split into 3 columns: the left column contains item from the given AUT problem as well as instructions, the middle column contains a text input box at the top for entering alternate uses, then a list of all uses written so far, and the right column contains the respective AI assistance.
@@ 3.1.1. Experimental conditions and treatment
Our experimental design involved two main phases: an exposure (paired) phase during which participants attempted the task with an LLM partner, and a test (solo) phase during which participants attempted the task on their own. Participants were given two-minute time frames per object to submit their ideas one by one, and could freely edit or delete any previously submitted ideas. We control for time per task to emulate real-world scenarios where knowledge workers will have the same deadline for a task, irrespective of whether they are using LLM or not. During the exposure phase, participants are introduced to three objects, one after the other. Every participant was randomly assigned to receive one of three types of LLM responses:
@@
The List of Strategies condition was designed to capture a common real-world use of LLMs: obtaining structured thinking or brainstorming frameworks rather than direct answers. Although this approach may introduce additional cognitive load during the exposure phase, it can potentially foster deeper, more enduring benefits observed during the test phase—an effect analogous to the long-term learning advantages reported in educational studies leveraging LLMs (Kumar et al., 2023; Bastani et al., 2024). The assigned type of LLM response appeared 5 seconds after the item was shown to the participant, and appeared character by character, similar to other chat LLM interfaces. Figure 2 shows the schema of the experiment design. Following the exposure phase, participants engaged in a brief distractor task to simulate forgetting, playing a game of Snake for one minute. In this subsequent phase, participants were assigned the Alternate Uses Task for a new object selected at random, this time without LLM support. None of the participants in any condition had LLM support in the test round. This was done to measure the residual effects of the use of LLMs in previous exposure rounds.
@@ 3.1.3. Dependent Variables
[…]
Figure 4 contains plots of the Alternate Uses Task ideas for the various divergent thinking dimensions (Segmented by phase and/or LLM response type). The left figure shows originality score from 2.68 to 2.80 on the Y-axis, and phase(Exposure and Test) on the X-axis. No LLM response shows an increase from 2.77 to 2.78 from Exposure to Test, LLM Answer shows an increase from 2.73 to 2.75, and LLM Guidance shows a steep decrease from 2.77 to 2.69. The middle figure shows the Average Number of Ideas from 4.6 to 6.0 on the Y axis and the phase(Exposure and Test) on the X axis. No LLM Response shows an increase from 5.5 to 5.7 across the phases, LLM Answer shows a slight increase from 5.2 to 5.4, and LLM Guidance has a steep increase from 4.7 to 5.6. Error bars for all data points are roughly 0.5 wide. The right figure shows Creative Flexibility from 53.5 to 56.0 on the Y axis and Condition (No LLM Response, LLM Answer, LLM Guidance) on the X Axis. No LLM Response and LLM Answer are 55.7 and LLM Guidance is 54.1, with error bars being roughly 1 unit long.
Figure 5. Plots of the average individual- (left) and group-level (right) median diversity, segmented by experiment conditions and phases. Higher values denote more difference between ideas. Error bars represent $\pm$ one standard error of mean.
[…]
Figure 5 shows two side-by-side plots of the average individual- and group-level median diversity, segmented by experiment conditions and phases. Higher values denote more difference between ideas. The average individual-level median diversity plot has Pairwise Cosine Distance from 37 to 40.5 on the Y axis. No LLM Response shows a slight increase from 39.4 to 39.7, LLM Answer shows a decrease from 40.0 to 38.8, and LLM Guidance shows a decrease from 38.0 to 37.3 (all from exposure phase to test phase). The average group-level median diversity plot has Pairwise Cosine Distance from 41 to 44 on the Y axis. No LLM Response shows an increase from 41.5 to 42.8, LLM Answer shows an increase from 42.4 to 43.4, and LLM Guidance shows a decrease from 41.7 to 41.2 (all from exposure phase to test phase).
@@ 3.2.1. Originality
[…]
Figure 6 shows plots of subjective measures collected before and after participants completed the Alternate Uses Task all segmented by the three LLM response types. The left figure shows participants’ change in self perceived creativity levels (Based on how many % of humans they felt they were more creative than) from -8 to 0, No LLM Response at -5, LLM Answer at -6, and LLM Guidance at -7, all with error bars roughly 5 units long. The middle figure indicates how their feelings towards the increased use of AI computer programs in daily life changed from 0 to 0.2, No LLM Response at 0.04, LLM Answer at 0.12, and LLM Guidance at 0.06, all with error bars roughly 0.16 units long. The right figure presents how much difficulty they had in coming up with ideas for the test object from 3.25 to 3.6, with No LLM Response at 3.53, LLM Answer at 3.33, and LLM Guidance at 3.51, all with error bars roughly 0.17 units long.
@@ 3.2.3. Creative Flexibility
[…]
Figure 7 illustrates the high level structural overview of the convergent thinking experiment. 640 participants start with a verbal fluency task, consent form, and pre-experiment survey before being randomized into one of three conditions for three exposure rounds. These conditions are: A (human-only), B (AI provides potential answers), and C (AI provides guidance). After the exposure rounds, participants engage in a 1 minute snake game as a brief interlude before moving onto 2 test rounds that all participants carry out without any AI support. Finally, they end with a feedback survey.
@@ 4. Experiment 2: Convergent Thinking
The results from our first experiment highlight how LLMs shape our ability to generate ideas independently. However, creativity involves more than idea generation; it requires the capacity to identify and refine the most appropriate solutions within a given conceptual space. Convergent thinking is crucial for advancing from a broad set of ideas to selecting the most viable outcome. Motivated by this, our second pre-registered [^2] https://aspredicted.org/4q75-bskd.pdf experiment focuses on convergent thinking, testing not only LLM-generated answers but also the potential of LLMs to guide users toward solutions, simulating a coaching approach.
@@ 4.1. Experimental Design
[…]
Figure 8 shows the 4 user interfaces used in the convergent thinking experiment, (No LLM Response, LLM Answer, and LLM Guidance exposure conditions, as well as for the test round). The No LLM Response and Test round interfaces are the exact same. A display showing the current round at the top left, and a timer at the top right. Right below this is some brief instruction, and right below that is space for participants to enter their answer. The LLM Answer and LLM Guidance interfaces have an extra section situated at the bottom on top of all the above. That is where the text containing the LLM assistance is being displayed.
@@ 4.1.1. Conditions and treatment
Similar to Experiment-1, the LLM Guidance condition was inspired by a common real-world method of using LLMs to obtain structured thinking frameworks rather than direct answers. After the exposure phase, participants played a game of Snake for one minute which served as a distractor task. In the test phase, all participants completed two additional RAT rounds (randomly selected) to attempt without any LLM assistance.
@@
During the exposure rounds, participants were made to wait the full 1 minute before advancing to the next task, while in the test rounds, participants could advance once they submitted an answer. This was to discourage participants from speeding through the experiment without giving thought to their answers. Like Experiment-1, we control for time per question to map onto real-world scenarios of LLM use in work environments. Participants were never shown the correct answer to the RATs they completed (except as part of LLM responses, if applicable), nor were they informed if their answers were correct.
[…]
Figure 9 shows Accuracy(%) from 50 to 80 on the Y axis, and phase (exposure and test) on the X axis. No LLM response shows a slight increase from 60 to 62, LLM Answer shows a decrease from 80 to 57 and LLM Guidance shows a slight decrease from 54 to 50.
@@ 4.1.2. Stimulus
For the Remote Associates Test (RAT), we employed a widely-used dataset from Bowden and Jung-Beeman (Bowden and Jung-Beeman, 2003). The subset we selected contained 45 questions equally distributed across three difficulty levels: easy, medium, and hard. For our study, we randomly selected questions from the easy and medium difficulty levels for each round, as pilot testing indicated that these were appropriate for our participant pool (crowdworkers). Figure LABEL:fig:rat-responses provides examples of the LLM responses presented to participants in the different experimental conditions.
@@ 4.1.5. Participants
[…]
Figure 10 shows 4 plots detailing participant perceptions across conditions, None, LLM Answer and LLM Guidance. The top left plot shows Change in Reported Self-Creativity Rating from -7.5 to 2.5 on the Y axis and condition on the X axis. No LLM Response is at -2, LLM Answer is at –3.5 and LLM Guidance is at -6. The top right plot shows Change in Feeling about AI Use from -0.10 to 0.10 on the Y axis and condition on the X axis. No LLM response is at -0.02 LLM Answer is 0.025 and LLM Guidance is -0.05. The bottom right plot shows Perceived Difficulty of Test Rounds from 2.4 to 2.7 on the Y axis and Condition on the X axis. No LLM response is at 2.52, LLM Answer is at 2.45 and LLM Guidance is at 2.66. The bottom right plot shows Percentage of Participants Who Found Exposure Rounds Helpful from 77.5 to 90 on the Y axis, and condition on the X axis. No LLM response is at 88 LLM Answer is at 83.5, and LLM Guidance is at 81.
@@ 5.1.2. Differential Impact of LLMs on Divergent and Convergent Thinking
Our experiments reveal that the effects of LLMs vary significantly depending on the aspect of creativity being measured. In divergent thinking tasks, where participants were asked to generate a wide range of ideas, we observed more skepticism toward LLM assistance. Participants seemed less inclined to adopt AI-generated suggestions, which may be due to the nature of divergent thinking itself—encouraging exploration and unconventional approaches (Runco, 1991; Runco and Acar, 2012; Baer, 2014; Guilford, 1967). In contrast, for the convergent thinking task, where participants were tasked with narrowing down the ideas to a single solution, LLM assistance during the exposure phase appeared more beneficial. This is consistent with the idea that convergent thinking is more structured and goal-oriented, making it easier for participants to recognize when LLMs are effectively guiding them toward the correct solution (Mednick and Halpern, 1968; Cropley, 2006; Simonton, 2015).
LLMs, however, may also hinder the creative process. In divergent thinking, the introduction of AI-generated ideas may distract participants, consume valuable cognitive resources, or prevent full engagement with the task. This aligns with theories of creativity such as Boden’s concept of conceptual spaces, which emphasize the importance of users exploring and navigating creative possibilities independently (Boden, 2010). Over-reliance on LLMs during divergent thinking may disrupt this exploration, leading to less original ideas, as seen in the originality results (Figure 4). Similarly, in convergent thinking tasks, LLMs might steer participants toward specific solutions too quickly, reducing the need for deep engagement with the problem space (Guilford, 1967).
@@ 5.1.3. Persistent Homogenization of Ideas and the Challenges of Designing LLM Coaches
Existing research has shown that LLM usage can lead to the homogenization of ideas within groups, where participants tend to converge on similar outcomes when using AI-generated suggestions (Anderson et al., 2024; Padmakumar and He, 2023; Doshi and Hauser, 2024). Our findings extend this concern, showing that even when people stop using LLMs that provide strategic frameworks for thinking, the homogenization effect can persist. In our divergent thinking experiment, participants who received LLM-generated strategies exhibited reduced diversity in their idea sets both during and after LLM use, suggesting that such frameworks may have a lasting impact on creative processes, potentially stifling the generation of more varied or unconventional ideas.
Interestingly, we did not find the same effect for participants who received direct ideas from a standard LLM. The List of Ideas condition did not lead to the same lasting homogenization once the LLM was no longer present. This indicates that while LLM-generated strategies can have long-term effects on creative diversity, providing ideas without a guiding framework may allow for more cognitive flexibility once participants stop using AI. Further, participants with no LLM interactions in the study collectively generated the most diverse idea sets (Figure 5). In our convergent thinking experiment, participants who received LLM guidance during exposure performed worse in the test round compared to those who received direct answers. This highlights the complexity of designing LLMs as coaches or guides. While direct answers may not always seem ideal, in some cases, they can be more effective than offering frameworks or strategies. Designers of Human-AI systems must take these findings into account, ensuring that LLM interactions are structured to avoid unintentional long-term effects on cognitive diversity. Controlled experiments, such as those conducted here, can be useful methods in refining and optimizing the design of coach-like LLM systems (Spatharioti et al., 2023; Kumar et al., 2023).
@@ 5.2. Broader Implications for Fields Involving Human-AI Co-Creativity
These insights have broad implications for various fields, from scientific discovery to the arts and humanities. In the context of AI for science, for instance, there is ongoing excitement about building increasingly sophisticated models to accelerate the process of hypothesis generation and experimental execution (Boiko et al., 2023; Si et al., 2024; Lu et al., 2024b). However, the design of these models often overlooks their potential impact on scientists’ creative abilities. Although some initial explorations have been conducted, the field lacks rigorous empirical evaluation of how AI systems affect human creativity in scientific discovery. The arts and humanities may face similar challenges (Rane and Choudhary, 2024). For example, writers using the same LLM could produce homogenized content, even when AI is no longer part of their workflow. This underscores the need for AI systems that not only assist in the creative process but also promote long-term cognitive diversity, ensuring that human creativity thrives in collaboration with AI rather than becoming constrained by it.
Our work in a controlled, simple setting suggests that the performance of AI models alone is not enough to determine their value. The design of human-AI interactions—whether AI provides direct answers or encourages users to think critically—and the timing of performance evaluation (during AI use versus after) both significantly alter the narrative. These considerations are critical in ensuring that AI models enhance, rather than diminish, human creative potential. Beyond simply building superhuman AI, we must focus on how these tools influence human creativity, culture, and cognitive growth, aiming for AI systems that enrich and elevate human thought (Brinkmann et al., 2023; Heersmink, 2024).